OpenStudy (anonymous):

The octane rating of a gasoline is a measure of the amount of isooctane in the gas. Subaru recommends 91-octane gasoline for 2008 Legacy 3.0 R. How much 87-octane gas and 93-octane gas should Kelsey mix in order to make 12 gal of 91-octane gas for her legacy?

OpenStudy (queelius):

Let x = gallons of 87 octane gas Let y = gallons of 91 octane gas Next, you need to set up some relationship to express the constraints. Hint: cx + ky / (c + k) is a weighted average that can be used in this problem.

OpenStudy (queelius):

Since you are logged off, I'll proceed with the next step. x + y = 12 We know that together, x + y must equal 12 gals. Next, we need to think about a mixture that gets us to 91 octane. To do this, we'll use the weighted average mentioned earlier. (87x + 93y)/(x + y) = 91 87x + 93y = 91x + 91y -4x + 2y = 0 So: x + y = 12 -4x + 2y = 0 Two equations, two unknowns. Solve. 6y = 48 -> y = 8 x = 4 So, 4 gallons of 87 octane gas, and 8 gallons of 93 octane gas will get us the desired octane.